an object of mass 3.0 kg moving with an initial velocity of 5.0i m/s collide and sticks to an object of mass 2.0 kg with an initial velocity of -3.0j m/s find the final velocity of the composite object.

1
Expert's answer
2022-05-09T14:49:48-0400

"u=\frac{\sqrt{m_1^2v_1^2+m_2^2v_2^2}}{m_1+m_2}=3.2~\frac ms."